The diameter of a round dinner plate is 10 inches. What is the area of the plate?
a- 314 square inches
b- 39.25 square inches
c- 78.5 square inches
d- 39.25 inches

Answer by Earlsdon(6098)

The area of a circle is given by: (pi)(r^2)
The given diameter is 10 inches.
The radius is half the diameter. r = 10/2 = 5 inches.
A = (3.14)(5^2)
A = (3.14)(25)
A = 78.5 square inches.